OpenStudy (anonymous):

If g(x) = 9 + x + e^x, find g−1(10).

OpenStudy (anonymous):

g-1= inverse of g(x)

OpenStudy (anonymous):

ok so first i get x=9+y+e^y

OpenStudy (anonymous):

then lnx=ln9+lny+y

OpenStudy (anonymous):

Im not sure if this is the correct method because I am unsure what to do next

myininaya (myininaya):

you are thinking way to hard

myininaya (myininaya):

you don't need to find the inverse function

myininaya (myininaya):

\[g^{-1}(10)=x \\ g(x)=10\] just look for an x such that 9+x+e^(x) will give you 10

myininaya (myininaya):

also taking ln( ) of both sides would not give you ln(x)=ln(9)+ln(y)+y it would give you ln(x)=ln(9+y+e^y)

myininaya (myininaya):

But anyways on this questions you are not actually expected to find the inverse function

myininaya (myininaya):

just the value

OpenStudy (anonymous):

ah i see thanks!

OpenStudy (anonymous):

ok so how do I solve for x from x+e^x=1

myininaya (myininaya):

you don't solve

myininaya (myininaya):

you think of a number x that will make 9+x+e^x , 10

OpenStudy (anonymous):

oh so 0

myininaya (myininaya):

yes

OpenStudy (anonymous):

thank u!

myininaya (myininaya):

\[9+0+e^0=9+0+1=10 \\ g(0)=10 => g^{-1}(10)=0\]

myininaya (myininaya):

np